Need help:
Evaluate the expression.
10C4
A) 210
B) 720
C) 2880
D) 3

Answer by checkley77(12844)

10C4
10!/4!*6!=(10*9*8*7*6*5*4*3*2*1)/4*3*2*1*6*5*4*3*2*1
10*9*8*7/4*3*2*1=5040/24=210 ans.